OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write a system of two equations with the variables m and f in each to represent this problem. pls dont goggle, answer it from what you know :) thx

OpenStudy (diyadiya):

m+5f=25 10f=40 (membership fee should only be paid once) therefore f=40/10=4 m+5(4)=25 m=25-20 m=5